A particle of mass is in a concentrated potential field V (𝑟). Determine the equation of motion using the Lagrangian! Is there a constant of motion? Explain!


1
Expert's answer
2021-07-05T17:45:54-0400

Given,

Mass of the particle = m

Potential field =V(r)

"L=T-V"

"L=m(\dot{r})^2- V(r)"


Now, "\frac{d}{dt}(\frac{\partial L }{\partial \dot{r}})=\frac{\partial L}{\partial r}"


"m\ddot{r}=\frac{\partial V({r})}{\partial r}"